# Combinatorics and Hodge theory of degenerations of abelian varieties: A survey of the Mumford construction

### v2

@@ 6. The second Voronoi fan and Alexeev’s theorem
We have a classifying map $Y^{\ast}\to{\mathcal{A}}_{g}$ and by the hypothesis on monodromy, we have, étale-locally about $0\in Y$, an extension and lift $Y\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}_{\underline{R}}}$, ${\mathbb{B}}_{\underline{R}}\coloneqq{\mathbb{R}}_{\geq 0}\{B_{1},\dots,B_{k}\}$, e.g. because this lift exists analytically-locally about $0\in Y$ (cf. Prop. 2.27). The morphism $Y\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}_{\underline{R}}}$ is algebraic, for instance by Borel algebraicity or by compactifying and applying GAGA. Pulling back $X^{\rm univ}({\underline{R}},\mathscr{H})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}_{\underline{R}}}$, which is algebraic and relatively projective by Proposition 6.12, we deduce the result. ∎
@@
By Proposition 4.10, there is a smooth quasiprojective variety $Y$, snc divisor $D\subset Y$, zero-stratum $0\in D$, and family $f^{*}\colon X^{*}\to Y^{*}$ of PPAVs over $Y^{*}=Y\setminus D$, whose monodromies about the components $D_{i}\ni 0$ are given by $r_{i}{\bf x}_{i}^{2}$. Applying Theorem 7.1, we deduce the existence of an extension $f\colon X\to Y$ with the desired properties.∎

### v3

@@ 6.2. Algebraicity and projectivity
We have a classifying map $Y^{\ast}\to{\mathcal{A}}_{g}$ and by the hypothesis on monodromy, we have, étale-locally about $0\in Y$, an extension and lift $Y\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}_{\underline{R}}}$, ${\mathbb{B}}_{\underline{R}}\coloneqq{\mathbb{R}}_{\geq 0}\{B_{1},\dots,B_{k}\}$, e.g. because this lift exists analytically-locally about $0\in Y$ (cf. Prop. 2.27). The morphism $Y\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}_{\underline{R}}}$ is algebraic, for instance by Borel algebraicity. Pulling back the family $X^{\rm univ}({\underline{R}},\mathscr{H})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}_{\underline{R}}}$, which is algebraic and étale-locally projective by Proposition 6.12, we deduce the result (the nodality and relative $K$-triviality of the extension follow as in the proof above). ∎
@@
By Proposition 4.10, there is a smooth quasiprojective variety $Y$, snc divisor $D\subset Y$, zero-stratum $0\in D$, and family $f^{*}\colon X^{*}\to Y^{*}$ of PPAVs over $Y^{*}=Y\setminus D$, whose monodromies about the components $D_{i}\ni 0$ are given by $r_{i}{\bf x}_{i}^{2}$. Applying Theorem 7.1 and passing to an étale chart about 0, we produce a projective extension $f\colon X\to Y$ with the desired properties.∎